OpenStudy (anonymous):

Given A(5,-8) and B(-6,2), find the point on segment ab that is 3/4 of the way from a to b

OpenStudy (anonymous):

Can anyone show the work for this one?

OpenStudy (anonymous):

\[(x,y) ≡ \left(\frac{x_1 + \lambda x_2}{1 + \lambda}\right),\left(\frac{y_1 + \lambda y_2}{1 + \lambda}\right) \] \(\lambda :1\) is the ratio in which the point divides the segment internally and the ratio in this case is \(\large\frac{\frac{3}{4}}{\frac{1}{4}} :1\). Find the point, using this and if the question is asking Length of the segment you can use the distance formula for it but you will need to tell the point from where you have taken (measured) the distance.

OpenStudy (anonymous):

Thanks for answering this for me it was a while ago but nice just the same.